OpenStudy (anonymous):

Solve the following system using the substitution method: -9x+3y=-15 y=3x-5

OpenStudy (anonymous):

-9x+3(3x-5)=-15 Is x=-1/3 or -3?

OpenStudy (anonymous):

-9x+3(3x-5)=-15 -9x+9x-15=-15 -15=-15 true x=any real number

OpenStudy (anonymous):

Ooooh ok, I see my mistake So, what's x so I can finish the problem?

OpenStudy (anonymous):

x can be any number; given some x, y is 3x-5 (0,-5) works, (1,-2) works, etc

OpenStudy (anonymous):

Oh ok, got it! Thanks!

OpenStudy (anonymous):

:)